OpenStudy (anonymous):

A sodium ion combines with a chlorine ion to form a molecule of sodium chloride, which is common table salt. A sodium ion has a net charge of +1 and a chlorine ion has a net charge of -1. What is the overall charge of 12 sodium chloride molecules

OpenStudy (unklerhaukus):

\[\text{Na}^+ + \text{Cl}^-\to\text{Na}^+\text{Cl}^-\to\text{NaCl}\]

OpenStudy (wmj259):

If you are combining the ions, basically elements with special charges, then on the left side you would have the individual ions with their respective charges and on the right side you would have the same amount of molecules as the left side had, ALSO with the same amount of charge that the left side had before. Which in this case you charges are canceled because a positive and a negative charge neutralize

OpenStudy (wmj259):

Basically, what @UnkleRhaukus said.

OpenStudy (unklerhaukus):

there is half as many molecules as there were ions,

OpenStudy (unklerhaukus):

\[12\text{NaCl}\leftarrow12\text{Na}^++12\text{Cl}^{-}\] \[Q{(12\text{NaCl})}= Q{(12\text{Na}^+)}+Q{(12\text{Cl}^{-})}\]

OpenStudy (unklerhaukus):

\[Q(12\text{NaCl})=12\times(+1)+12\times(-1)=\]

OpenStudy (anonymous):

sorry i didn't see it until now thanks